Mr. Ali has a job as a pharmacist in Shaheen chemist earning PKR-30,000 per year, and he is deciding whether to take another job at Walt chemist as the manager of for PKR-40,000 per year or to purchase a franchise of Watson chemist that generate revenue of PKR-200,000 per year. To purchase Watson chemist Ali would have to use PKR-20,000 savings and borrow another PKR- 80,000 at an interest rate of 10 percent per year. The running expenses of new pharmacy would be PKR-80,000 for supplies, PKR-40,000 for labor, PKR-10,000 for rent, and PKR-5,000 for utilities. Assuming zero tax and repayment of principal loan after three years, answer following questions. a) What would be business and economic profit of Jones in case he purchase Watson chemist? b) If there will be a new pharmacy opens up in close vicinity after three years, what would be economic profit of the Watson chemist in three years? c) What theory of profit will apply on the three years earning of Watson chemist?
1
Expert's answer
2017-09-27T10:40:09-0400
Salary 1 = PKR30,000 per year,
salary 2 = PKR40,000 per year,
revenue = PKR200,000 per year,
savings = PKR20,000,
borrowing = PKR 80,000 at interest rate of 10 percent per year,
expenses: PKR80,000 for supplies, PKR40,000 for labor, PKR10,000 for rent, and PKR5,000 for utilities.
a) If Jones purchase Watson chemist, then:
business profit = 200,000 - (80,000 + 40,000 + 10,000 + 5,000) = PKR65,000.
economic profit = 65,000 - (30,000 + 8,000) = PKR27,000.
b) If there will be a new pharmacy opens up in close vicinity after three years, then economic profit of the Watson chemist would be lower in three years.
c) The theory of economic profit will apply on the three years earning of Watson chemist.